[ a field is in the form of a parallelogram , whose perimeter is 450 m and one of its sides is larger than the other by 75 m find the length of all sides ]
NOTE: it's not clear whether you meant to use meters or centimeters (you gave perimeter in meters and side in cm), but it's one or the other, so I'm assuming its meters.

Answer by TimothyLamb(4379)

---
2x + 2y = 450
x = y + 75
---
put the system of linear equations into standard form
---
2x + 2y = 450
x - y = 75
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= long sides = 150 m
y= short sides = 75 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php